# Argument-Aware Approach To Event Linking

### v1

@@ Appendix C Implementation Details
For the analysis of negative data generation methods, details on data generation are covered Appendix B. Additionally, the KB pruning baseline is implemented by randomly pruning 10% of the unique labels present in the training set. We then label the corresponding samples (5984 samples) as out-of-KB.
For the LLM baseline, since existing event linking systems do not use LLMs, we take inspiration from existing methods Pradeep et al. (2023) to apply an instructed LLM (GPT-3.5-Turbo) to re-rank the candidates retrieved by the bi-encoder in a zero-shot fashion. Due to the cost of GPT-3.5-Turbo, we sample 1000 test examples and use the same retrieved candidates for both our method and the LLM baseline to re-rank. For the Wikipedia test set, the model is asked to simply re-rank the top-k documents given the query. For the NYT test set, the model is given an additional option to simply label the query as “NIL” if none of the documents describes the event in the query. The two prompts for the Wikipedia and NYT test sets are shown in Figure 6 and Figure 7, respectively.
Appendix D Full Bi-encoder Evaluation Results

### v2

@@ Appendix D Implementation Details
For the analysis of negative data generation methods, details on data generation are covered Appendix C. Additionally, the KB pruning baseline is implemented by randomly pruning 10% of the unique labels present in the training set. We then label the corresponding samples (5984 samples) as out-of-KB.
For the LLM baseline, on the Wikipedia test set, the model is asked to re-rank the top-k documents given the query. On the NYT test set, the model is given an additional option to simply label the query as “NIL” if none of the documents describes the event in the query. The two prompts for the Wikipedia and NYT test sets are shown in Figure 6 and Figure 7, respectively.
Appendix E Full Bi-encoder Evaluation Results